OpenStudy (anonymous):

Find the monthly payment of the given loan. A loan of 10000 dollars amortized over 48 months at 8%

OpenStudy (ar43r):

275 10000*0.08*4=3200 10000+3200 = 13200 - total sum to payoff 13200/48=275

OpenStudy (anonymous):

hmmmm that wasn't the correct answer

OpenStudy (ar43r):

what is the correct answer?

OpenStudy (anonymous):

i don't know! i get three tries and that last answer was incorrect

OpenStudy (ar43r):

amortized means payoff? 8% per year?

jimthompson5910 (jim_thompson5910):

monthly payment is 244.13

OpenStudy (ar43r):

how you calculate it?

jimthompson5910 (jim_thompson5910):

Use the formula \[M=\frac{Pr/12(1+r/12)^n}{(1+r/12)^n-1}\] So I got (10000*(0.08/12)(1+0.08/12)^48)/((1+0.08/12)^48-1) = 244.13

OpenStudy (ar43r):

thanx. never saw something like this. can you tell how this formula is called, so I can look on the wiki?

OpenStudy (anonymous):

Derivation Of mortgage formula P=initial balance R= interest rate W=monthly payment \[1 \text{st} \text{Month}\text{ }= P(1+R) -W\] \[2 \text{nd} \text{Month}\text{ }=((P(1+R) -W)(1+R))-W\] \[P(1+R)^2-W(1+R)-W\] \[3\text{rd} \text{Month}=\left(\left(P(1+R)^2-W(1+R)-W\right)(1+R)\right)-W\] \[P(1+R)^3-W(1+R)^2-W(1+R)-W\] \[N \text{th} \text{Month}= P(1+R)^N-W\left(1+(1+R)+(1+R)^2+\text{...}.(1+R)^{N-1}\right)\] \[\text{Using} \space \text{Geometric} \space\text{Sum} \text{Formula}\] \[\left(1+(1+R)+(1+R)^2+\text{...}.(1+R)^{N-1}\right)=\frac{\left(1-(1+R)^N\right)}{1-(1+R)}=-\frac{1-(1+R)^N}{R}\] \[P(1+R)^N-W\left(-\frac{1-(1+R)^N}{R}\right)\]\[=P(1+R)^N+W\frac{\left(1-(1+R)^N\right) }{R}\]

OpenStudy (anonymous):

W= monthly payment R= Monthly payment P= initial \[10000\left(1+\frac{.08}{12}\right)^{48}+W\frac{\left(1-\left(1+\frac{.08}{12}\right)^{48}\right) }{\frac{.08}{12}}\text{==}0\] W=244